### v1

# Exploiting Kubernetes’ Image Pull Implementation to Deny Node Availability

Abstract
Kubernetes (K8s) has grown in popularity over the past few years to become the de-facto standard for container orchestration in cloud-native environments. While research is not new to topics such as containerization and access control security, the Application Programming Interface (API) interactions between K8s and its runtime interfaces have not been studied thoroughly. In particular, the CRI-API is responsible for abstracting the container runtime, managing the creation and lifecycle of containers along with the downloads of the respective images. However, this decoupling of concerns and the abstraction of the container runtime renders K8s unaware of the status of the downloading process of the container images, obstructing the monitoring of the resources allocated to such process. In this paper, we discuss how this lack of status information can be exploited as a Denial of Service attack in a K8s cluster. We show that such attacks can generate up to 95% average CPU usage, prevent downloading new container images, and increase I/O and network usage for a potentially unlimited amount of time. Finally, we propose two possible mitigation strategies: one, implemented as a stopgap solution, and another, requiring more radical architectural changes in the relationship between K8s and the CRI-API.
Index Terms:
Kubernetes, Container Security, Cloud Computing, eBPF.
I Introduction
Since its inception, Kubernetes (K8s) has been conceived as a container management system to orchestrate multiple workloads in a scalable, resilient, and reliable way [1]. This containerization deployment-oriented approach enables cluster and application automation, multi-tenancy, and reproducibility. Additionally, K8s’ loosely coupled architecture based on asynchronous communication, isolation, and decoupling of concerns pioneered a remarkable change in the cloud-native computing landscape.
Being implemented with a modular structure, K8s uses a set of APIs to manage the network, storage, and container runtimes. This creates a scenario where several modules with distinct implementation and requirements need to communicate to execute a task (e.g., the creation of containers), usually in an asynchronous way. Such a design provides high availability, scalability, and performance but creates gaps that could potentially be used as attack vectors.
In this paper, we investigate the hazards of the asynchronous communications between two modules inside the K8s cluster, the Kubelet, the module that runs on each node and communicates with the central API, and the container runtime, which is tasked with handling the creation, management, and deletion of containers in K8s nodes. Our study figures that since the communication between these modules happened asynchronously, it does not maintain any intermediate status on each request, enabling a malicious user to exploit this behavior as a threat that can be used as a vector to several distinct attacks.
In particular, when deploying a Pod— a group of one or more containers — an image pull is usually triggered for the corresponding containers if unavailable locally. If a container is deleted before the image is fully downloaded, the API will not communicate this action to the container runtime. However, the runtime will still download the image, wasting resources in the process. Similarly, when a container is forcibly destroyed, its resources are not instantly released. While K8s will delete all information from the API as soon it receives the request, the complete elimination of the containers composing the Pod will only occur when all modules finish their deletion processes. Although this is intended behavior, such a process could be exploited to mount a Denial of Service (DoS) attack on the cluster. Indeed, these image download requests would prevent other users from scheduling new jobs without a cached image and force the tenant to consume unnecessary resources.
To understand the actual impact, we present several possible attack scenarios, studying the impact on the scheduling, system resources, and other containers running in the node. We also describe in detail the threat, with its limitations and requirements. We show how this issue affects all Kubernetes deployments, demonstrating its effect both in private and public cloud settings. Additionally, we discuss how this could be solved in K8s. Since the solution is not trivial and will require time and effort to be achieved, we present a temporary, eBPF-based proof-of-concept solution to address the problem until the fix is fully implemented in all the relevant projects.
We also argue that the majority of K8s security research over the last ten years focused on lateral container movement, container escapes, and authentication and access control[2, 3, 4, 5, 6, 7]. These topics have been extensively researched, and several ideas have been proposed to address these issues. However, studying the interactions between K8s and its runtime interfaces is one area that is frequently disregarded both in the literature and in practice, mainly due to the challenges presented by a modular project with several distinct working groups. Taking that into account, the contributions of this work are summarized as follows.
New resource-based DoS threat in K8s. We demonstrate that it is possible to exploit a series of design flaws in the Kubernetes CRI-API to execute a DoS attack. We also discuss the importance of API status routes and how the asynchronous communication between distinct modules can be used as a vulnerability point.
Attacking strategies. We explore how such an attack can overcome a cluster’s defenses and mount a DoS attack on a target node, blocking the image download of other applications and hampering the resource use by the containers. We also show that it is possible to manipulate the node’s image cache, increasing the time needed to start applications in the cluster.
eBPF proof-of-concept solution. We discuss how the problem can be mitigated, with the pros and cons of each solution. We thoroughly explain how the current design flaws in K8s’s CRI-API could be efficiently fixed. As a stopgap solution, we propose a proof-of-concept that employs the extended Berkeley Packet Filter (eBPF) to intercept and mitigate rogue downloading images.
II Background
This section provides the technical background for the remainder of this work. The following subsections present a small review of the Kubernetes, the containerized application lifecycle, and the Container Runtime Interface.
II-A Kubernetes
Kubernetes is an open-source orchestration system for automating deployment, scaling, and management of containerized applications, initially developed by Google and maintained by the Cloud Native Computing Foundation (CNCF) [8]. K8s is built from a set of composable modules through a set of standard APIs that can be extended by the users alongside the core components [1]. Over time, this allowed the development of several new infrastructure paradigms, like Function as a Service (FaaS) frameworks[9] and multi-cluster management [6].
K8s uses a modular structure to implement all its services, each with a distinct function, in the cluster. The API Server is the central communication hub for the entirety of K8s. It processes requests, validates them, and maintains the desired cluster state. After processing the request, the API Server is responsible for forwarding the desired state to other modules, like the Kube Controller or the Kube Scheduler. Furthermore, the API Server enforces access controls and authentication mechanisms, ensuring secure and managed access to cluster resources.
To deploy a containerized application on K8s, the minimum manageable object is the Pod. A Pod is a group of one or more containers that share resources, shown as a single access point for the other objects in the cluster. Pods are designed to be ephemeral, and represent a single instance of an executable resource in K8s. Rarely created as an individual object, it usually uses a more complex controller, such as StatefulSets and Deployments, to manage its deployment and replication on the cluster nodes.
II-B Container Runtime Interface
K8s relies on a well-defined API called Container Runtime Interface, also called CRI-API or just CRI, which communicates through gRPC Remote Procedure Calls[10] with the container runtime to instantiate the Pods in the nodes. Today, the default runtime used by K8s is containerd, also maintained by the CNCF. The Kubelet module, that is the node agent for the K8s API, uses the CRI to execute the operations in the container runtime. Finally, the container registry serves as a centralized repository for storing, organizing, and distributing container images to nodes when required. The registry additionally enables a secure and efficient way to manage the entire lifecycle of container applications.
The Container Runtime Interface (CRI) was developed as a replacement for Dockershim, which previously served as the bridge between K8s and Docker. The introduction of CRI aimed to standardize communication with container runtimes, eliminating the necessity for intermediary solutions like Docker. This standardization enhances compatibility and fosters a more modular approach, allowing seamless integration with emerging container technologies such as Kata Containers and gVisor [11].
II-C Container Image Deployment
To better understand how the Pod lifecycle is managed in the cluster, Figure 1 shows the sequence diagram for three kubectl commands: first, when a Pod deployment is initiated using the run command; second, when that same Pod is deleted; and finally, when the deletion is forced with the ---force option. In this last scenario, the API does not wait for the CRI to confirm the deletion status. It instead immediately removes any related entries in etcd (K8s’s database) once the command is entered.
Figure 1: Pod Creation and Deletion sequence diagrams
We can break the execution of all commands into two distinct steps. Initially, the user interacts with the API that redirects the command to the correct node. The second step is executed on the selected node asynchronously with respect to the first step. Only at the end the status of the execution is returned to the API. While, by default, the instantiation happens asynchronously, the deletion will wait until a response is returned. To change this behavior, kubectl implements two options, ---force and ---grace-period=0, that execute the deletion without waiting for the Pod completion process.
Implemented to improve the agility of the cluster management, these options can, as described in the K8s official documentation [^1] https://kubernetes.io/docs/concepts/workloads/pods/pod-lifecycle/, be disruptive for some workloads, or some resources may continue to run indefinitely on the cluster. We exploit this behavior to show that resources could be used by the node without any notice, mainly by how the CRI was developed.
III Motivation and Attack Requirements
In this section, we first present the threat model and assumptions of our work. Second, we describe our motivation. Afterwards, we describe the requirements to execute the attack on K8s, including the configuration and image parameters that can increase the attack surface and its impact.
III-A Threat Model and Assumptions
The attacker aims to increase resource usage inside a node without being noticed by the administrator. That behavior can be used as a single DoS attack or an auxiliary action to more complex threats. We assume that the attacker has access to a Service Account (SA) or an user credential from a compromised Kubelet config file, or through an exploited container with a SA linked to them. This SA requires a minimal set of permissions for the attack: creation and deletion of Pods.
As described in the Role-Based Access Control (RBAC) Best Practices official documentation [^2] https://kubernetes.io/docs/concepts/security/rbac-good-practices/, it is expected that a single cluster will have several distinct SAs to limit the access to specific namespaces, to share the cluster between production and development, or to enable a continuous integration/delivery solution. So, it is reasonable to assume that one of these service accounts could be leaked from a development machine, a non-protect staging deployment used as a vector to access an inside container with a mount SA token, or a insider malicious user with access to the cluster.
We also assume that quotas and limitations on the number of API requests per second are active in the cluster. Further security configurations, e.g., image validation and locked-down private registries, can harden the attack execution but not wholly mitigate it. Indeed, a well-tuned configuration can make the threat viable even in a very constrained scenario.
On the restriction side, we understand that a compromised container or a SA token could be used for more advanced attacks, such as lateral movement [7], Out-of-Memory (OOM) attack [12], or deploying cryptomining in the node [13]. However, all these attacks can be straightforwardly identified by inspecting the resource usage or checking the elements running in the cluster. Notwithstanding, this attack wants to create a stealthy channel to cripple the cluster’s resources that cannot be easily identifiable without an extensive evaluation of the node resources and processes, differing it from the other ones.
On the other hand, the attack can be performed with the same minimal permissions on a wide variety of settings, ranging from busy clusters with several concurrent deployments to smaller, soft multitenancy clusters [^3] Soft-multitenancy refer to resource sharing and strong trust between the tenants. https://kubernetes.io/docs/concepts/security/multi-tenancy/#isolation, and multi-cluster topologies [14]. Even with monitoring or auditing logs active, we argue that the attack will be hard to discover without a specific and deeper analysis of the node itself, which is not a trivial task. Indeed, even in cloud deployments — such as Google Kubernetes Engine (GKE), Azure Kubernetes Service (AKS), and Amazon Elastic Kubernetes Service (EKS) — monitoring tools tend to provide an aggregated view of the cluster’s resource usage, and administrators usually will not access directly the cluster nodes.
We further assume that the node and guest Pods have no known vulnerabilities and that all security mechanisms work properly.
III-B Asynchronous API Communication Status
As described in subsection II-C, K8s uses an asynchronous communication to enforce the actions in the several interfaces implemented in the cluster. For example, each Pod creation is executed independently by K8s, but each node’s containerd will have the last say in defining the order of execution for each request.
This behavior can cause problems since it is necessary to control the operations’ atomicity to guarantee the consistency of the cluster. Furthermore, some operations are part of a bigger dependency tree that must be executed in a certain order. For instance, deploying a container requires the download of its corresponding image to be completed (if not already available).
Since after sending the request K8s maintains no control over the request execution, it relies on the interface implementation to update the intermediate and final status on the object. However, this does not always happen. The CRI has no API related to the status of image download, and so, the only response containerd will provide will be either the download confirmation or some error on the downloading process, like a missing secret or an unavailable tag.
The Kubernetes Enhancement Proposals (KEP) 3542444GitHub issue available at https://github.com/kubernetes/enhancements/issues/3542 proposes an update to the Kubelet, enabling download status updates to the API, solving the problem previously described. Even with this KEP, the problem is not entirely solved: indeed, it remains impossible to cancel pending downloads.
III-C Motivating Example
The attacker aims to select one or more worker nodes within the Kubernetes cluster and starve their resources, eventually causing a resource DoS condition. Once the attack has been successfully mounted, nodes can no longer schedule new Pods whose images are not already present in the local cache. Besides that, the CPU, disk, and network usage will also dramatically increase.
To attain this, attackers can exploit the aforementioned lack of status updating in CRI for image downloads. The attack strategy is simple: the attacker sends a Pod creation request, waits for the API to register that the image needs to be downloaded, and then deletes the Pod. This will not interrupt the download that has already been started. Repeating this multiple times will cause a substantial, sustained increase in resource consumption.
As said before, this attack is hard to identify and mitigate, since the requests are legitimate, and could be easily confused with a normal user making some API requests. On the other hand, the node will suffer from an high and possibly irregular workload, as each image is downloaded and unpacked. Furthermore, the inability to download images may be attributed to an external issue with the container registry itself, and not to a local DoS attack.
IV Attack Description
The following section proposes two attacks that exploit the aforementioned vulnerability in containerd and the CRI-API, first rendering the Kubernetes node unable to pull any Pod image for an extended period and, secondly, impacting a node’s cache, increasing the amount of time needed to deploy an instance of an application.
IV-A Attack Vector
As mentioned in subsection III-B, KubernetesAPI abstracts away several steps that must be taken behind the scenes to create a Pod. When a user creates a Pod through the API (as in Figure 1), it communicates with the Kubelet to issue a Pod creation request. The Kubelet then communicates with containerd via the CRI to trigger the image pull, set up the container, and finally the Pod status is returned to the user. In normal conditions, the API waits for containerd to respond with the status of the Pod creation before sending a response to the user. However, the API does not wait for containerd to complete pulling the images. In fact, as soon as all required containers for a Pod have been created by containerd, the API immediately sends a successful response to the user.
While this choice has its roots in the separation of concerns between Kubernetes and containerd, it has a subtle side effect. Without having to notify the API about the status of the image pull, containerd is completely unaware of the addition or removal of Pods through the Kubernetes API. With that, a malicious actor can try to use a significant amount of resources by repeatedly instantiating new Pods in the node. However, K8s addresses this with the implementation of Quotas, which limit the number of objects that can be created in a namespace. The malicious actor can overcome this limitation by deleting objects immediately after the instantiation. Nevertheless, Quotas will still apply until the object is fully removed from the cluster.
As discussed in subsection II-C, one of the ways to immediately release resources is to use the ---force flag. This will automatically remove the API bindings for the resources. On the other hand, it will not interrupt any image download that has already been started by containerd. [^5] At the time of writing, a bug exists in K8s that additionally prevents the deletion of associated containers when a Pod is force deleted. See https://github.com/kubernetes/kubernetes/issues/119276. Thus, after deploying a Pod an attacker can force delete it shortly after, essentially bypassing Quotas and avoiding having to wait for the image downloads to complete. This process can be repeated a potentially infinite number of times, in quick succession. By doing so, an attacker can clutter the containerd’s download queue with useless image pull requests as long as they have image tags available to pull. Since containerd is agnostic to the Pods’ existence and Kubernetes is unaware about what images are being downloaded, the status quo can remain indefinitely. As such, the target container runtime will continuously be forced to keep track of and download useless container images on behalf of the attacker.
IV-A1 Attack Requirements
The attack requires minimal prerequisites to be executed. First, it needs access to a cluster with standard user privileges and the capacity to create and delete at least one Pod. Any limitation imposed by the RBAC configuration, Quotas, and Pod Security Standards is irrelevant to its execution. Second, the attacker needs a set of Docker images, with different tags of a single image, or entirely different ones altogether. They also need to be in a registry accessible by the node. Furthermore, the least layers the images have in common, the better, since re-occurring layers will not be re-downloaded.
The content, structure, functionality, and size of the image are irrelevant for the attack to be successful. Their only limiting aspect is that the total size of the image cannot exceed eventual limits imposed by the cluster administrator because the download will be blocked before starting.
On the other hand, the amount of images available to the attacker is critical for the attack’s success. Since the image (or any of its layers) must not be present on the attacked node, the more are available, the longer the attack will last. This must also scale up with the amount of storage available on the node.
IV-B Denial of Resources
This attack will generate a denial of resource in the node, and its impact is twofold. First, the endless cycle of image pull requests and deletions prevents the deployment of new containers without a cached image. While this impacts less long-running tasks such as web services, it prevents short-lived workloads, such as Continuous Integration/Continuous Delivery (CI/CD) pipelines, from deploying and completing in a few seconds or minutes, causing a loss of availability. This problem is exacerbated by the growing trend of running short-lived microservices: indeed, 72% of containers live fewer than five minutes [15]. An unresponsive cluster can be extremely problematic, even if the attack is short lived: clusters that rely on short and fast batch jobs (e.g. CI/CD) will be inevitably delayed.
Second, the attack drains the available resources on the worker node by forcing it to constantly utilize disk, bandwidth, and CPU to retrieve container images. As a result, other workloads hosted on the node may suffer a performance setback, possibly violating Service Level Agreements. Since the Kubelet usually runs as a process in the node and it is not directly controlled by the quotas and limitation imposed by the Kubernetes, it will use 100% of the resources available to download and decompress the images, and by default, there is no way to limit this resource access.
IV-C Cache Manipulation
Since the images are not deleted after the download, they can be used to manipulate the cache in a given node. The Garbage Collector (GC)’s default configuration sets a series of thresholds, the most important ones being ImageGCHighThresholdPercent, used to determine when the GC will be triggered, and ImageGCLowThresholdPercent, the minimum target percentage that the GC wants to achieve after starting to delete images. The default settings for these two thresholds are 85% and 80%, respectively.
After surpassing the upper threshold, the Kubelet marks images present on disk for more than two minutes (by default) as eligible for deletion, and starts removing them until at least the lower threshold is reached. It must be noted that the GC may keep deleting eligible images even after surpassing the lower threshold. This can be a problem in scenarios where the cluster hosts cron jobs, FaaS tasks or any application that does not have a container continuously running. Since their corresponding images will be deleted, such tasks will suffer from a delayed execution, as containerd will be forced to download the image again.
However, K8s does not just limit to cleaning up unused images and containers. Another key part of the system is the Node-pressure Pod eviction system [^6] https://kubernetes.io/docs/concepts/scheduling-eviction/node-pressure-eviction/, which is independent from the GC. When a node’s resources reach dangerously low thresholds, the Kubelet intentionally terminates Pods to avoid possible contention and crashes. At the time of writing, hard eviction is handled by the DefaultEvictionHard option, whose default value is 90%. Once the threshold is reached, it will delete both Pods and their corresponding images, taking care in deleting as few of them as possible. This is meant to be an extreme measure, and usually causes further disruption on the node due to the critical lack of disk space and increased CPU usage.
IV-D Responsible Disclosure
The attack presented in this work has been responsibly disclosed to the Kubernetes Security Team. They informed us that DoS via Pod creation is considered a resource usage bug. Nevertheless, we are actively collaborating with them to address this issue, and we have opened a public issue  [^7] The link for the issue to solve this in future versions.
V Experiments and Results
In this section, we present a summary of our findings after running a series of experiments. We evaluate the effectiveness of the attack on a real K8s cluster and additionally analyze the effect of the attack on some key system resources.
V-A Experimental Setup
We set up a local testbed on our premises to evaluate the attack. On a machine equipped with an Intel Xeon Silver 4112 CPU (2.60GHz), we created a virtualized environment with a master and two worker nodes without any special software configuration. Each machine has 2 vCPUs, 4 GB of memory, 120 GB of storage capacity (HDD), a 1 Gbit/s network interface, and Ubuntu 22.04 LTS as its Operating System (OS).
While to the best of our knowledge no information about the average size of K8s nodes can be found in literature, we tried to match the specifications of our nodes to the default one that users find when creating a K8s cluster on GKE [^8] GKE uses a e2-medium machine type with 2 shared vCPUs, 4 GB of memory and 100 GB of storage capacity per node., Linode999www.linode.com/pricing/ or DigitalOcean101010www.digitalocean.com/pricing/kubernetes. On the other hand, AKS does not provide default values but suggests slightly higher ones in their best practices [^11] https://github.com/Azure/k8s-best-practices/blob/master/Cost_Optimization.md#node---vm-sizes.
Finally, we used a separate physical machine to create an additional Virtual Machine (VM) acting as a private Docker registry. The registry and the cluster also communicate through a 1 Gbit/s network interface.
To further confirm our findings, we additionally deploy a cluster on GKE, using n2d-standard-2 machines with 2 AMD Epyc vCPUs, 8 GB of RAM, a 100 GB balanced persistent disk and a 10 Gbit/s network interface. All the experiments outlined in this section have been successfully replicated in both the public and private cloud scenarios, with comparable results. Due to space constraints and unless specifically mentioned, the following results and graphs refer to the private cloud experiments.
V-A1 Images
For the effectiveness and consistency of the experiment, we decided to build our own set of Docker images to be used in the attack.
By researching the Docker image compression mechanism, which relies on gzip, we discover that different types of randomly generated data influence the decompression time of archives. For example, random English text is efficiently and quickly compressed by gzip, while random data formatted with the od command yields varying times and compressing rates. The details of this research are available on GitHub [16].
Exploiting these results, we create two sets of images to assess the attack’s effectiveness on two distinct image sizes: one with large layers focused on bandwidth usage and the second with small layers focused on the I/O overhead.
Each image in the set contains a base Ubuntu layer (of around 80 MB uncompressed), which is shared between all the images, and a variable number of layers, created with the od command and the c option. Moreover, by forcing the --no-cache flag, we make sure that every layer is always re-built and, therefore, unique in terms of content.
The first set, called Variable GB, contains a total of seven images. Each of the images contains one to seven additional layers of 2GB each. As a result, each image in the set ranges in size between 2GB and 14GB. The second set of images called Variable MB, contains forty images, with each additional layer containing 20 MB of data. This results in images ranging between 100 MB and 900 MB. These sizes refer to on-disk usage, as when fetched from a registry, they will be compressed, decreasing their size by roughly half. Table I shows a summary of these two sets.
Image set Variable GB Variable MB
Tag count 7 40
Lower-bound uncompressed size 2.12GB 0.10GB
Upper-bound uncompressed size 14.40GB 0.90GB
Total uncompressed size* 57.12GB 16.80GB
Total compressed size* 28.88GB 8.47GB
TABLE I: Summary of layers, size, and other characteristics of the two image sets used in the experiments. *The total size includes the base image only once, as it is not duplicated when generating an image. The compressed size takes into account pigz’s compression factor of 0.504.
V-B Denial of Resources and Scheduling Delay
(a) CPU usage by type with $mp$ set to 1 (top), to 2 (middle) and to 4 (bottom).
(b) Disk write and network read with $mp$ set to 1 (top), to 2 (middle) and to 4 (bottom).
Figure 2: CPU usage by type (on the left); disk write and network read (on the right) of the system during the variable GB scenario. Data were taken from a randomly sampled trial.
We focus on assessing the denial of resources capabilities of the attack. In particular, we investigate to what extent the Pod scheduling can be delayed and how the CPU and I/O are affected.
V-B1 Scheduling Delay Metric
To properly compare the different lengths of the attack and the actual impact on the node image pulling, we created a metric that would normalize the results. We call it Scheduling Delay. It is measured in s/GB and is calculated as follows:
$SD=\frac{D}{S_{c}}$
Where $D$ is the duration of the attack (in seconds), and $S_{c}$ is the total amount of compressed data that the node had to withstand from the beginning of the attack. This metric is unrelated to the impact on system resources and gives us a rough idea of how efficient is a certain attack configuration in preventing a node’s downloading of new Pod images.
V-B2 Max Parallel Image Pulls
Starting from version v1.27, K8s allows users to specify the maximum number of images to be pulled in parallel [^12] GitHub issue available at https://github.com/kubernetes/enhancements/issues/3673; indeed, in the past, images were exclusively pulled serially. While this feature helps improve the efficiency of the cluster by making full use of the available bandwidth, we want to investigate its effects on the attack since it would prevent the cluster from being stuck in a given download. However, this option would probably increase resource consumption.
In the experiment, we use the values 1, 2, and 4 of the MaxParallelImagePulls flag. This flag will be shortened to $mp$ in the upcoming sections for brevity.
V-B3 Results
Variable GB set. We first run the experiments of the Variable GB image set. Before each experiment, we reshuffle the images. Then, for each image, we deploy it, wait two seconds, and then delete it with the ---force flag. Additionally, to verify the attack’s behavior with the parallelism provided by the $mp$ flag, we run each experiment with the values described before. We repeat this process thirty times for each setting; the final result is an average of these runs. To better visualize the results, we chose a random trial from the thirty and plotted the CPU usage, broken down by type, and the disk write and network read. These plots are available in 1(a) and 1(b).
With $mp$ set to 1, the results are staggering. Even with a small attack, no parallelism and just seven images, the attack generates a significant impact on the cluster, which clearly can be categorized as a DoS. Every new download on the node is blocked for a total of 1352 seconds, almost 23 minutes; the CPU usage averages 65% with spikes of almost 90%, and the disk is kept constantly busy, with I/O write spikes of 150 MB/s. From 1(a) (row 1), we can see how these CPU peaks are noticeable and correlate with the download of the images (1(b)), and thus with a high bandwidth and I/O usage. On the other hand, when images are decompressed, the CPU usage drops significantly but remains steadier, as does the disk usage.
As we increase $mp$, the total time of the attack decreases, and the resource usage increases accordingly. For example, setting $mp$ to $2$ reduces the attack time to 909 seconds, while with $mp$ $=4$, it further reduces to 887 seconds. On the other hand, the CPU usage increases, while the difference in the peaks becomes less and less apparent as the various parallel downloads compete for CPU and I/O time with the ongoing unpacking tasks. With four parallel images, the CPU usage reaches almost 100% for most of the attack time.
By calculating the Scheduling Delay for these three tests, we see how it also drops from 46.82 ($mp$ $=1$) to 31.48 ($mp$ $=2$) and 30.72 ($mp$ $=4$) s/GB. It is important to observe the difference between the former two settings, where the SD drops 33%, and the latter, where the SD drops again but by a smaller margin (only 2.4%). While introducing parallelism strengthens the attack and puts an increased load on resources, at a certain point, the returns start diminishing as contention is introduced, and the cluster’s overall performance starts degrading. Thus, we believe that increasing $mp$ without implementing a solution to this threat could be even more detrimental.
Variable MB set. We now repeat the same experiments on the Variable MB set. Such results are similar to the ones we obtained with the Variable GB set, but with a much lower runtime due to the overall smaller size. Thus, we calculate the Scheduling Delay metric, presented in Table II, and the average CPU usage over all cores, presented in Table III, to properly assess the consequences of using smaller-sized images in the attack.
Set Scheduling Delay
mp$=1$ mp$=2$ mp$=4$
Var. GB $46.82\pm 0.37$ $31.48\pm 1.02$ $30.72\pm 0.72$
Var. MB $56.93\pm 1.59$ $41.10\pm 1.16$ $39.92\pm 1.60$
TABLE II: Scheduling Delay (s/GB) averaged over 30 trials with standard deviation. The higher, the more disruptive the attack.
Set CPU Usage
mp$=1$ mp$=2$ mp$=4$
Var. GB $67.31\pm 0.81$ $91.67\pm 2.24$ $93.22\pm 2.23$
Var. MB $64.14\pm 0.35$ $84.35\pm 0.62$ $85.78\pm 2.69$
TABLE III: Average CPU usage over all cores of a worker node during the attack for each scenario. Factor of 100, averaged over 30 trials.
From the Scheduling Delay results, the immediate impression is that using smaller images is indeed more disruptive. With $mp$ $=1$, the Variable GB set hangs the cluster for 47 seconds per GB. On the other hand, the Variable MB manages to block it for 57 seconds. Thus, we can block the downloads for 18% more time with small images. With $mp$ $=2$ and $mp$ $=4$, this becomes more evident, arriving at 23% more with small images. We attribute this increase to the overall overhead that containerd must withstand when downloading images. Indeed, each new image implies fetching its manifest, obtaining the list of layers, and manually downloading each of them: all operations repeated many times start piling up.
On the other hand, Table III presents almost opposite results. While with no parallelism, the difference is almost unnoticeable, as $mp$ is increased, the Variable GB set is far more CPU-intensive, using almost 10% more CPU on average than the Variable MB set. We attribute this lower average CPU usage to the relative weight of downloading and unpacking a smaller image. As a consequence, the possibility of saturating the CPU is lower unless the parallelism is way higher than in our experiments.
Finally, we repeat our experiments in the GKE deployment as outlined in subsection V-A, obtaining consistent results. As GKE has yet to introduce support for the MaxParallelImagePulls parameter, we only test the scenario with sequential downloads. The Scheduling Delay of the attack decreases due to the increased bandwidth available in the GKE cluster. Additionally, a slightly weaker CPU contributes to an overall increase in CPU usage during the attack. However, such slight differences do not change the validity of our previous observations.
V-C Service Interference
Benchmark No attack With the attack
Results Trials Results Trials
kernel $823.25\pm 0.41\>s$ 3 $1630.04\pm 5.47\>s$ 5
stress $2185.83\pm 1.29\>bo/s$ 3 $1463.18\pm 12.52\>bo/s$ 15
TABLE IV: Summary of the results of the benchmark on the non-disrupted and on the attacked cluster. $bo/s$ stands for bogo-operations/seconds.
Another goal of the attack is to affect the Quality of Service (QoS) of other running services in the node. We perform two experiments with the same attack, but this time, we verify how this affects CPU and I/O usage of other Pods.
To do so, we employ the Phoronix Test Suite[17], an automated benchmark suite that allows users to perform a wide range of performance assessments on their system. From the wide variety of benchmarks available, we choose the build-linux-kernel [^13] Specifications available at https://openbenchmarking.org/test/pts/build-linux-kernel-1.15.0. The “defconfig” option was used. and stress-ng [^14] Specifications available at https://openbenchmarking.org/test/pts/stress-ng. The “cpustress” option was used. benchmarks, two of the most popular of the suite. The first assesses the time needed to compile the Linux kernel (version 6.1) in the machine, putting an I/O and CPU workload on the system. The latter evaluates the amount of operations per second the CPU can perform. To ensure the statistical relevance of its results, Phoronix automatically performs additional trials if it encounters performance issues or suspiciously high variability between different runs.
First, we run the build-linux-kernel benchmark with no attack running on the cluster. In this case, it completes in 853.25 seconds, averaged over three runs. Then, we apply the attack as described in subsection V-B with the Variable GB set and $mp$ set to 1 (as parallelism is not a concern in this setting). We choose this particular scenario as it is the longest and, thus, the most likely to last the entirety of the benchmark. We therefore re-run it, applying the attack in the background and cleaning up after every trial. With this setup, the build time increases to 1630.04, a 91% increase over the non-disrupted one.
The other benchmark also reports significant results, with the stress-ng measuring 2185.83 bogo-ops/s [^15] Bogus operations per seconds, a measure used by stress-ng to evaluate processor performance with no attack running over three trials. Repeating it with the attack, Phoronix takes 15 trials to obtain 1463.18 bogo-ops/s, a drop of 33% with regards to the non-attacked scenario. Additionally, the standard error is 12.52 seconds, almost ten times higher than the 1.29 seconds of the non-attacked benchmark. This 33% drop is less severe than the one from build-linux-kernel, mainly due to the fact that stress-ng’s CPU-bound behavior clashes more with the attack, introducing additional contention.
To further confirm our findings, we replicate our experiment in the GKE public cloud scenario, in identical conditions. With this setup, the resource impact is still significant. The build-linux-kernel benchmark over 5 trials increases the build time by 80%, while stress-ng reports an average of 45% less bogo ops/s over 15 trials. Thus, the impact on the benchmarks run on a public cloud is evident and comparable to the one on private cloud. Since on this setup the CPU is slightly weaker and the bandwidth is greater, the images are downloaded faster, but the CPU is used more in the process. Thus, CPU-bound benchmarks such as stress-ng suffer a greater drop than on the original environment. On the other hand, the disk on the GKE VM is faster, resulting in a lesser impact on the build-linux-kernel benchmark.
V-D Cache Manipulation
Figure 3: Sequence diagram of downloaded images (left axis) and used disk space (right axis, shaded area) during the sequential image download. Dark lines represent active downloads; grey lines represent image locality (i.e. image finished downloading and present on disk).
Finally, we verify the attack’s capabilities in interfering with another kind of system resource: storage. In particular, we analyze how the attack can manipulate the local image cache.
As previously described in subsection IV-C, K8s is equipped with a local image cache which is periodically expunged by a GC should the used disk space surpass the ImageGCHighThresholdPercent (85%) threshold. Thus, we aim to investigate if the attack can have the side effect of inducing cache evictions. Such scenario would mean that the attacker can have a way of indirectly controlling the cache. In addition, it would further disrupt the node’s operations as the I/O and CPU overhead of the cache eviction is added to the underlying attack.
We therefore run again the Variable GB experiment with $mp$ being set to 1, as we are not interested in seeing the effects of parallel downloads in this case. With a simple script, we monitor the images present in the cache and the disk space. However, this time, we run the attack sequentially rather than randomly to eliminate variability. In other words, first we deploy the smallest image (2 GB), delete it with ---force, deploy the second smallest, and so on. We repeat the experiment several times, obtaining varying results. An example result is shown in Figure 3.
Indeed, we discovered that the GC is extremely aggressive, and once the 85% threshold has been reached (at $t=13$), it starts deleting all qualifying images that are eligible for deletion. In our case, during the download of the seventh image the GC deleted images numbered 1 to 4 at once, dropping the disk usage to 62%. Since the minimum Time to Live (TTL) of images downloaded to disk is two minutes, the sixth image was not removed. On the other hand, the fifth image was spared for no reason, and in other trials, it was also deleted.
Either way, this greedy behavior by the GC, while beneficial for the disk pressure, further bolsters the capabilities of the attack. Since in the background more images are being downloaded, saved to disk and unpacked, triggering the deletion of several images at once causes further impact on system resources and potentially deletes images that may have been needed for short jobs by other tenants.
On the other hand, we obtained a slightly different result by avoiding to delete Pods at all. Doing so prevents the GC from deleting any image — as they are all in use — and forces it to wait for the eviction manager to delete at least one Pod. Consequently, the disk usage eventually reaches 90%, triggering the eviction manager. As it evicts Pods, the GC also deletes their image shortly after. This overlap of both the eviction and GC causes even more mayhem on the node, heavily impacting CPU and storage I/O.
V-E Attack Feasibility
The results presented in the subsubsection V-B3 and subsection V-D are strongly related to the resources available in the node. That opens an important discussion point about the attack feasibility generalization. We understand that generally, K8s worker nodes will have a limited size, mainly because of the applications’ nature running on it — as they are stateless — which usually do not maintain any status. In addition, bigger nodes will commonly be linked with large applications. That incentivizes a more spread use of resources between distinct nodes (i.e., horizontal node autoscaling) instead of a resource increase in the nodes already deployed.
These scenarios are ideal for small resource-based DoS attacks in specific nodes, where a continuous flow of CPU and I/O resources can deny a given application or force the eviction to another node. The same could be pernicious in Edge nodes, where the constrained resources with no redundancy are highly vulnerable to this kind of attack. Finally, since the spread usage of the resources and expected applications’ burst usage, it is almost impossible, with the current tools, to mitigate the threat only with resource monitoring.
Figure 4: Architecture of the MAGI System. Dark components are part of K8s; light components are part of MAGI. Arrows represent communications and data flow.
VI Mitigation Discussion
This section presents the rationale about how the problem can be solved in the Kubernetes, containerd, and CRI-API code, the difficulties in implementing the solution, and, as an alternative, we present a proof-of-concept solution, using eBPF until all the necessary implementation has been released inside the respective projects.
VI-A Kubernetes Status and Best Approach
To some extent, the limitations of the CRI-API in controlling/reporting the image downloading process are already known by the community. As mentioned in subsection III-B, KEP-3542 has already been assigned to implement an image progress notification. Open since 2019, the reasoning behind this proposed enhancement is to bring more information to the administrator, reporting the current speed and download status from each image. Nonetheless, the threat previously described will not be solved by this implementation since the progress notification will not enable the termination of the downloading process if a container has been marked for deletion.
On the other hand, KEP-3673 outlines the Parallel Image Pulls option (outlined in subsubsection V-B2) and its role in preventing cluster saturation due to prolonged image downloads. We show in our experiments how in reality this feature can actually increase the effectiveness of the attack: with more images to be downloaded concurrently, the attack will consume more resources.
As discussed in subsection III-B, the problem is the lack of communication status routes in the CRI-API, because there is no status message between the user’s request (create or update) and its completion. The difficulties to completely solve this issue are threefold. First, ensure that the containerd maintains the context open during all the command executions. Second, implement a series of new API routes to guarantee that the CRI can cancel the submission or the download of each requested image to the runtime. Finally, the Pod controller must also be adapted to include these new routes, maintaining the status of the requests, validating if the image download has already started, and deciding if it needs to stop after a delete request.
Unfortunately, these implementations are not trivial and depend on several interactions with different Special Interest Groups (SIGs) and projects with distinct schedules and contribution methodologies. We are already collaborating with them to solve these issues. However, the implementation process is usually very long and laborious. Therefore, we developed a temporary mitigation solution that is presented in the following subsection.
VI-B MAGI System
As a stopgap and proof-of-concept solution, we propose the MAGI System (Mitigation AGainst Images) [18], whose architecture is shown in Figure 4. It employs eBPF, Python, and Bash and monitors the cluster status and image downloading requests, taking prompt action whether an attack matching our pattern is detected.
The MAGI System is composed of two main components: the master component and the node component. The master component, as the name suggests, must run on the master node of the cluster; node components must be run on each node of the cluster. This includes the master itself, if users wish to schedule Pods on it.
Master component
The master component leverages Kubernetes’s auditing capabilities to keep track of the Pod creation and deletion requests, which are handled by the API server. As a pre-requisite, thus, it requires enabling Kubernetes’s Auditing feature [^16] https://kubernetes.io/docs/tasks/debug/debug-cluster/audit/, which by default is disabled. The Auditing features lets cluster administrators specify with great flexibility what to monitor, for example user activity, API usage, and control plane events. For our purposes, we limited the scope and verbosity of the logging to only the pods/* resource group. More precisely, we captured all API calls related to the Pods group and subgroups, logging event metadata. In the case of elaborate topologies with multiple master nodes, it is sufficient to place the master component in a node in which the auditing capability has been enabled.
Once the Auditing is properly set up, the master component can be started. During regular usage, it will monitor the API server’s calls for Pod creations, keeping a cluster-wide track where each Pods have been scheduled and if any image pull is necessary for them. Once the image has been downloaded and the Pod has been successfully deployed, the component discards the collected information.
When the component instead detects a Pod deletion request, it checks if the Pod had previously requested an image pull which has yet to complete. Such an event could potentially signal an attack attempt. In this case, the master component promptly contacts the node in which the Pod was scheduled, informing it that the tracking image should not be downloaded anymore.
Node component
The node component is responsible for several concurrent tasks, and works with three threads. The first is in charge of keeping track of the images being downloaded by containerd; the second monitors the API calls being made by the Kubelet to containerd; the third listens for the alerts emitted by the master component, eventually performing corrective actions. Each instance of the node component operates independently from the others.
Image download queue monitoring. The first thread keeps track of crictl’s local image download queue. This information is fed to the script by containerdsnoop [19], a fork of dockersnoop [20], an eBPF program capable of intercepting gRPC calls to and from containerd; in particular, calls from the ImageService (crictl) and the RuntimeService (Kubelet).
Within the implementation, gRPC calls are not encrypted, but the HTTP2 HPACK header compression mechanism requires access to the initial packets of connections. Thus, before the startup, the Kubelet instance is deliberately restarted, allowing our tool to fully parse the calls. While the program can parse almost every API call made to containerd, we narrowed down its scope for our tool, deliberately snooping only requests and responses to the PullImage API route.
When the tool detects a request, the payload of the call is inspected and the requested image is appended to the queue. When a response is detected – signaling a successful download – the said image is removed from the queue. The tool is able to distinguish between Docker Hub and other-registries images, and maintains the ordering of the download queue even when the MaxParallelImagePulls parameter is set to a value greater than 1.
System call log monitoring. A second thread is tasked with monitoring the system calls being made by containerd. Once we know that a certain image is being downloaded, we also need to identify which connection is being used to perform the download of which layer. Unfortunately, containerd does not expose such information natively. However, we fortunately noticed that containerd’s behavior is predictable.
When containerd downloads an image, it first opens a connection to the registry, retrieving the manifest file and obtaining the list of layers that compose the image. Then, for each layer, containerd spawns a thread that creates a folder on the disk for storing the downloaded files, then immediately after binds to a network socket. By default, containerd opens a maximum of four sockets at a time [^17] This option is customizable within containerd, and is unrelated with the MaxParallelImagePulls K8s flag, which instead refers to parallel images rather than parallel layers. when downloading images.
To properly log this information, we used a custom-made eBPF-based tool [21] that specifically monitors two selected system calls. In particular, we attached eBPF keyprobes to the mkdirat and tcp_connect system calls being performed by containerd. The first is related to the creation of the folders containing the layer data on disk and the latter with the opening of the connections to download the data in question. Since the two steps are performed by the same containerd thread, by correlating the two calls we are able to identify the mappings between containerd threads, layers, and open ports. This effectively allows us to understand which connection is being employed to download which layer. The two tools continously fetch new data and push it to userspace using a perf_buffer.
Alert listening and actuator. Finally, a thread is tasked with listening for the master’s alerts. When such an alert arrives, the component promptly reviews its local image pull queue: if the requested image is currently being downloaded, the component terminates the connection between the node and the registry, interrupting the download and clearing the local cache.
This is done first by querying the bbolt database181818bbolt is a fast on-disk database used by containerd. and obtaining the list of layers from the name of the image. Alternatively, the component can also query the registry directly, obtaining the list of layers from the manifest file. From each layer in the list, the port number of the connection downloading that particular layer is obtained. Finally, each of the offending sockets is terminated using ss -K. Since the Kubernetes API has already deleted the Pod and freed up its resources, it will not attempt again to download the image, thwarting the malicious actor’s attack. Clearly, if the attacker were to try again to download said image (or any other one), the system would get triggered again, thus being able to continously take down attack attempts over time.
To account for pending downloads, we put offending images signalled by the master component in a blacklist. This behavior is triggered when an image download is requested but one or more images are in the queue before it. As the queue empties, the script automatically detects if a new download matches an image in the blacklist and terminates it accordingly.
VI-C MAGI Evaluation
To evaluate the effectiveness of our mitigation approach, we devised a series of three tests that evaluate the functionality of the system in different scenarios: without the attack, with the attack, and with the attack and the mitigation running.
Each scenario comprises the deployment of a series of images in a K8s cluster at pre-defined intervals. We selected some common images with different sizes, to represent a common cluster scenario. While in real clusters workload might be scheduled more irregularly and unpredictably, we devised a series of ordered commands with a constant wait time between each other, allowing us to accurately assess the behavior of the cluster and our solution while minimizing external interferences. Since the $mp$ parameter is not the focus of this experiment, we set it to 1.
In particular, in each scenario we perform the following steps. First, to emulate the deployment of a large-sized image, we deploy sagemathinc/cocalc in background and wait 20 seconds to give K8s sufficient time to complete the deployment and propagate the changes. Then, depending on the setting, we either start the attack (using the variable GB image set) or do nothing, and we wait 20 further seconds. Finally, to simulate the download of small images, we deploy jupyter/scipy-notebook, nginx, and online-boutique [22], waiting 20 seconds between each other. In particular, online-boutique is a collection of twelve lightweight microservices deployed as separate containers.
We repeat these steps for each of the aforementioned scenarios. In each one, we measure the total length of the test, the moment at which every deployment was requested and the moment at which the container was actually started. This information is collected using the in-built K8s APIs.
In the first scenario, the total time spent by the test is around eleven minutes. Of these, eight of them are spent downloading the cocalc image, which is extremely large; once the download is completed, the other images follow suit quickly. On the other hand, in the second scenario we deploy the attack after the first twenty seconds, thus queuing it after the cocalc download. Once cocalc finishes downloading, all the following requests are brought to a grinding halt as the attack starts blocking the cluster. This increases the length of the test to around twenty-nine minutes. It is important to note that the attack does not affect the net download time of the images; rather, it makes them stall and wait until their turn arrives. Finally, in the third scenario, the MAGI System successfully intercepts and blocks six out of seven images, i.e., all images but the smallest one. As the downloads get interrupted almost instantly, the total length of the test drops back to eleven minutes, the same as the first scenario.
These tests highlighted the capabilities of our solution. First, we observed how MAGI has almost no footprint. The CPU usage in nodes spikes at 0.8% during detection and almost 0% when sitting idle. The memory footprint was more evident, ranging from 380 MB idle to almost 420 MB during the attack, but remaining mostly stable nevertheless. We attribute this usage to Python’s memory management and the need of storing status for several different metrics at once. On the other hand, both CPU and memory usage were insignificant on the master node, with the only discernible overhead being K8s’ auditing feature.
However, the fact that MAGI failed to block the smallest image of the attack further strengthens the need for a more systematic and structured approach to the problem. While the eBPF code always identifies all offending downloads, the process of tracing back to the original layer is cumbersome. With enough bandwidth, small images get downloaded so quickly that it is almost impossible to even attempt blocking their download. In our opinion, the sheer amount of information required for gracefully terminating a download without disrupting any service, the inherent unpredictability of the system scheduler and the relative opacity of containerd’s download management render any further attempts in increasing MAGI’s speed better spent in improving the main problem instead.
VII Related Work
Denial of Service based on Resources Usage. Resource-based DoS in Cloud Computing was previously covered in previous works. In [23], the authors present a novel class of attacks called Resource-Freeing Attacks (RFAs) and create a methodology to prevent the DoS by increasing the amount of resources available to the victim VM. Fang et al. also discussed resource and co-location attacks in [24] and [25], where they proposed respectively one new metric, called Heteroscore and a threat method, REPTTACK and present a mitigation technique to this type of attack. Finally, in [26], Zhan et al. show a framework that aims to detect CPU-exhaustion DoS attacks in containers called Coda. However, it does not consider the node resources, just the container ones.
Kubernetes and Container Security. Ahmed [27] shows that the container deployment process can generate significant resource usage, mainly memory and CPU. The author also tested several alterations to the container deployment’s decompress process, focusing on the parallelism of several simultaneous layers. Such approach decreases the total time to deploy a given image but inherently increases the resources used in this process. Kubernetes is also focused in two papers on vulnerabilities in the configuration of distinct services in the orchestrator. He and Guo [3] show a threat in the resource configuration, where using highly privileged configurations allows eBPF applications to escape containers and access sensible information. Xiao [4], instead, found several threats based on the communication between microVMs solutions — such as Firecracker and Kata Containers — and containerd to escape attacks and resource-based DoS.
Another important area that has received recent attention from the container orchestration community is anomaly detection in cluster resource usage. Almaraz-Rivera [5]’s study highlights the importance of implementing monitoring and alerting systems to track containers’ performance and resource utilization. He and Guo also implement a demo solution using an ML auto-encoder algorithm. Lastly, Yolchuyev [28] proposes an eXtreme Gradient Boosting (XGBoost) based model for anomaly detection on the Kubernetes orchestration platform. Either case, to the best of our knowledge no prior work has been conducted on exploiting the asynchronous design of the K8s API communication.
VIII Conclusions
This paper demonstrates that the asynchronous implementation of some routes from the CRI-API can be exploited to create a resource-based DoS. We illustrate this threat through three distinct attacks. First, we focus on resource usage, where the attack used up to 95% of the CPU. Second, we show how this attack could interfere with other applications running in the cluster. Finally, we investigate how the cache could be manipulated by the constant stream of requests performed by the attack. We further experimentally prove the effectiveness of each attack in a local testbed and on GKE. Finally, we provide a guideline to solve the problem in all the related projects and are currently collaborating with the working groups to solve this issue as soon as possible. As a stopgap, we present an eBPF-based proof-of-concept to monitor the activity in the cluster, stopping useless image downloads.
Acknowledgments
This work was partially supported by project SERICS (PE00000014) under the MUR National Recovery and Resilience Plan funded by the European Union - NextGenerationEU and by European Union’s Horizon Europe Programme under grant agreement No 101070473 (project FLUIDOS).